# Self-Play Fine-Tuning Converts Weak Language Models to Strong Language Models

### v2

@@ 6.1 Experiment Setup
In this study, we adopt zephyr-7b-sft-full as our base model. This model derives from the pre-trained Mistral-7B (Jiang et al., 2023) and has been further fine-tuned on the SFT dataset Ultrachat200k [^1] https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k by HuggingFace. Ultrachat200k represents a high-quality 200k subset of the larger UltraChat (Ding et al., 2023) corpus, which comprises approximately 1.4M dialogues produced using OpenAI’s Turbo APIs. From UltraChat200k, We randomly sample $50$k prompts and use the base model to generate the synthetic responses. We subsequently follow the optimization method described in Section 4.1 for further training. In multiple iterations, we leverage the synthetic data from the most recent iteration and add to the newly generated synthetic data, therefore resulting in a synthetic dataset size of $50$k at iteration 00 and $100$k at iteration $1$, $2$ and $3$. At each iteration, we train our model for $2$ epochs.
Evaluation.
@@
We employed the widely used Huggingface Open LLM Leaderboard (Beeching et al., 2023) as our evaluation benchmark, using the same Language Model Evaluation Harness library (Gao et al., 2023b). This leaderboard encompasses 6 different datasets, each focusing on a a specific capability of LLMs. Collectively, these datasets provide a thorough assessment framework, evaluating LLMs on commonsense reasoning (Arc (Clark et al., 2018), HellaSwag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2021)), multi-task language understanding (MMLU(Hendrycks et al., 2020)), human falsehood mimic (TruthfulQA (Lin et al., 2021)) and math problem solving (GSM8k (Cobbe et al., 2021)). In evaluation, the language models are prompted with few-shot in-context examples and the question. We follow the standard approach and report the average score across all datasets. In Table 1, we detail the evaluation setting adopted by both the leaderboard and our experiments. We leave further implementation details to Appendix B.
@@
6.2 SPIN Effectively Improves Benchmark Performance
We demonstrate the effectiveness of SPIN using HuggingFace Open LLM Leaderboard as a wide range of evaluation. In Table 2, we compare the performance of our fine-tuned model by SPIN after iterations 0 to 3 with the base model zephyr-7b-sft-full. We can observe that SPIN exhibits remarkable effectiveness in improving the model’s performance by further leveraging the SFT dataset, on which the base model has already been fully fine-tuned. At iteration 00, where model responses are generated from zephyr-7b-sft-full, we observe an overall improvement of $2.66\%$ on the average score. The improvement is particularly significant on the TruthfulQA and GSM8k benchmarks, with improvement exceeding $5\%$ and $10\%$ respectively. At iteration $1$, we employ the LLM model from iteration 00 to generate new responses for SPIN, adhering to the procedure outlined in Algorithm 1. This iteration yields further enhancements of $1.32\%$ on average, and especially significant on the Arc Challenge and TruthfulQA benchmarks. Subsequent iterations continue this trend of incremental improvement across various tasks. Meanwhile, the improvement at iteration $t+1$ is naturally smaller than that at iteration $t$. As the iterative training progresses, the degree of improvement gradually approaches zero, suggesting that the model has reached a limiting point in the last iteration.
@@ 6.2 SPIN Effectively Improves Benchmark Performance
SPIN iteration 00 63.40 49.18 72.69 35.10 84.38 60.03 $60.80_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\pgfsys@color@rgb@stroke{0.0}{0.5}{0.0}\pgfsys@color@rgb@fill{0.0}{0.5}{0.0}(+2.66)}}$
SPIN iteration $1$ 65.19 55.17 72.30 35.78 84.96 59.34 $62.12_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\pgfsys@color@rgb@stroke{0.0}{0.5}{0.0}\pgfsys@color@rgb@fill{0.0}{0.5}{0.0}(+1.32)}}$
SPIN iteration $2$ 65.96 54.91 73.56 38.06 85.41 59.93 $62.97_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\pgfsys@color@rgb@stroke{0.0}{0.5}{0.0}\pgfsys@color@rgb@fill{0.0}{0.5}{0.0}(+0.85)}}$
SPIN iteration $3$ 65.87 54.90 73.72 38.97 85.54 59.99 $63.16_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\pgfsys@color@rgb@stroke{0.0}{0.5}{0.0}\pgfsys@color@rgb@fill{0.0}{0.5}{0.0}(+0.19)}}$
Comparison with DPO.
zephyr-7b-beta is a model derived from zephyr-7b-sft-full, trained with DPO on approximately $62$k preference data. This data, the UltraFeedback Binarized dataset(Cui et al., 2023) [^2] https://huggingface.co/datasets/HuggingFaceH4/ultrafeedback_binarized, comprises both chosen and rejected completions evaluated by GPT-4. We note that, DPO requires either human input or advanced language model feedback to determine the preference, making data generation a rather expensive procedure. In contrast, our SPIN only requires the initial model itself. Moreover, unlike DPO which requires new data source, our method exclusively leverages the existing SFT dataset. In Figure 3, we show the performance comparison of SPIN at iterations 0 and 1 (employing $50$k SFT data) with DPO training, from the same SFT checkpoint. We can observe that, while DPO leverages more data from new sources, SPIN based on the existing SFT data can already achieve comparable average performance to DPO training at iteration 0. From iteration 1, SPIN even surpasses the performance of DPO on the leaderboard benchmark.
6.3 Ablation Studies
In this subsection, we examine the effect of synthetic dataset size and training epochs within an iteration. Our analysis demonstrates the effectiveness of the synthetic data used by SPIN compared to the SFT data, as well as the necessity of iterative training in SPIN. Furthermore, to comprehensively assess the performance improvements of SPIN, we perform additional evaluations on benchmark tasks distinct from those in the Open LLM leaderboard.
Training Size.
We investigate the effect of varying training data size on the performance of SPIN. In Figure 4, we demonstrate the effect of training size for SPIN during iteration 00 and additionally compare with SFT with the full original dataset. Specifically, for the SFT baseline, we fully fine-tune Mistral-7B on Ultrachat200k for three epochs and report first epoch performance as the starting point (with x-axis 0) in the figure for SFT. For SPIN, we report the zephyr-7b-sft-full checkpoint as the starting point, which has also been fine-tuned on Ultrachat200k for one epoch. We select the training size of SPIN at iteration 0 to be 14k, 26k, and 50k and generate the data accordingly, ensuring that the larger dataset encompasses the smaller dataset. The performance of SPIN was then evaluated after 1 epoch of self-play fine-tuning for each training size. We can observe that, while SPIN results in notable improvement with increasing training sizes, SFT on further epochs 2 and 3 fails to yield more than $1\%$ improvement. Lastly, in Table 3, we also show the performance of SFT from zephyr-7b-sft-full on Ultrachat200k for one epoch. While self-play fine-tuning with synthetic data from zephyr-7b-sft-full effectively improves its performance, simply fine-tuning it again on the SFT data leads to degraded performance, as similarly observed in Figure 4.
@@ 6.3 Ablation Studies
Iterative Training v.s. Training for More Epochs.
We further study the training within iteration 00 and compare with the performance achieved in iteration $1$, particularly contrasting the test performance obtained from extended training duration with that from next iteration. Figure 5 depicts the performance trajectory of the model trained using SPIN over multiple epochs at iteration 0. It is evident that the most substantial improvement occurs during the first two epochs, followed by only modest gains in subsequent epochs. Notably, SPIN exhibits robustness and stability; extending the training duration does not diminish performance but rather maintains a rather consistent level. Nevertheless, the observation suggests an inherent limitation to the performance achievable within a single iteration, thereby underscoring the necessity for iterative training. As shown by the test performance achieved at iteration 1 in the figures, extending the training in iteration 0 fails to reach the performance comparable to iteration 1.
(a) Arc Challenge accuracy.
(b) TruthfulQA score.
(c) Average score.
Further Investigation on More Tasks.
Here, we further investigate the performance of SPIN on a broader variety of tasks, including MT-Bench (Zheng et al., 2023), Big-Bench (bench authors, 2023) and OpenBookQA (Mihaylov et al., 2018) in addition to the Open LLM Leaderboard tasks. Specifically, we use the following tasks from Big-Bench-Hard for a more comprehensive evaluation, including Causal Judgment (causal reasoning), Sports Understanding (commonsense reasoning) and Formal Fallacies (logical reasoning). In Table 4, we show the resulting scores of SPIN on MT-Bench as well as those tasks from Big-Bench. In Figure 6, we detail the model performances on MT-Bench with regard to different types of questions. We can see a notably robust improvement in the performance of SPIN on various tasks besides the HuggingFace Benchmark, without major degradation. Notably, on MT-Bench, the model fine-tuned by SPIN has surpassed the performance of vicuna-13b-v1.5 (Chiang et al., 2023) with a score of $6.57$.
@@
SPIN iteration 00 $6.46_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\pgfsys@color@rgb@stroke{0.0}{0.5}{0.0}\pgfsys@color@rgb@fill{0.0}{0.5}{0.0}(+0.52)}}$ 57.75 51.6 95.2 46.8
SPIN iteration $1$ $6.65_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\pgfsys@color@rgb@stroke{0.0}{0.5}{0.0}\pgfsys@color@rgb@fill{0.0}{0.5}{0.0}(+0.19)}}$ 58.82 51.2 95.2 47.2
SPIN iteration $2$ $6.78_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\pgfsys@color@rgb@stroke{0.0}{0.5}{0.0}\pgfsys@color@rgb@fill{0.0}{0.5}{0.0}(+0.13)}}$ 59.36 51.2 94.4 47.6
@@
7 Conclusion and Discussion
Limitation and Future Work.
Our theoretical results demonstrate that the optimization process of SPIN converges if and only if the LLM’s distribution aligns with $p_{\mathrm{data}}$. Therefore, our study focuses on a fixed target data distribution generated by humans, which inherently imposes a ceiling on the performance of fine-tuned LLM. Exploring the dynamically changing target data distribution is an important direction to overcome this limitation and elevate the LLM’s performance beyond this ceiling or even to a super-human level. Moreover, considering the resource demands of synthetic data generation, another promising avenue for further exploration is to reduce the volume of required synthetic data.
Appendix A Further Related Work
Curriculum Learning.
In deep learning, it has been observed that training models using data samples arranged in a strategically meaningful order can lead to improved performance compared to training on randomly shuffled data. This approach is commonly known as curriculum learning (Bengio et al., 2009; Soviany et al., 2022). Initial studies in curriculum learning introduced efficient algorithms that adhere to an ‘easy-to-hard’ progression (Spitkovsky et al., 2009; Kumar et al., 2010; Lee and Grauman, 2011; Zhang et al., 2015). In the field of Natural Language Processing (NLP), criteria such as sentence length and term frequency are commonly utilized (Cirik et al., 2016; Zhang et al., 2018; Liu et al., 2018). More recent developments include the application of curriculum learning algorithms in multi-modal learning (Liu et al., 2021; Wu et al., 2022). Our work shares a similar idea to curriculum learning, wherein the training data evolves iteratively—beginning with responses that are easy to distinguish from human-annotated data and gradually progressing to more challenging instances.
Generative Adversarial Learning.
Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) represent a distinct class of generative models, characterized by their unique adversarial process. To enhance training stability and data quality, Mao et al. (2017) introduced the Least Squares GAN, employing a least squares loss function for the discriminator. A significant advancement in GANs involves the use of Integral Probability Metrics (IPM) (Müller, 1997), particularly highlighted in the development of Wasserstein GAN by Arjovsky et al. (2017). This model employs IPM in its loss design, enhancing training stability. Since then, IPMs have become popular in the design of GANs (Mroueh and Sercu, 2017; Gulrajani et al., 2017), particularly in constraining the discriminator to a specific function class, thereby preventing it from overpowering the generator. Furthermore, Jolicoeur-Martineau (2018) generalized IPM-based GANs by introducing relativistic discriminator and proposed Relativistic GAN. It is worth noting that the objective function defined in our (4.2) is similar to Relativistic GAN (Jolicoeur-Martineau, 2018) and reduces to an IPM framework in Wasserstein GAN (Arjovsky et al., 2017) with a linear loss. However, our approach differs in both the choice of the function class and the training procedure. Inspired by GAN, Cheng et al. (2023) proposed an adversarial learning framework named Adversarial Preference Optimization (APO) that trains the LLM and a reward model in an adversarial game. Our method is also related to Generative Adversarial Imitation Learning (GAIL) (Ho and Ermon, 2016), which trains separate discriminator and policy networks in each iteration for imitation learning. In contrast to the above methods, SPIN relies on self-play where both the main player and the opponent player are the same LLM from two consecutive iterations.
Appendix B Experiment Details
B.1 Hyperparameters and Implementation Details

### v3

@@ 6.1 Experiment Setup
Model and Datasets. In this study, we adopt zephyr-7b-sft-full as our base model. This model derives from the pre-trained Mistral-7B (Jiang et al., 2023) and has been further fine-tuned on the SFT dataset Ultrachat200k [^1] https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k by HuggingFace. Ultrachat200k represents a high-quality 200k subset of the larger UltraChat (Ding et al., 2023) corpus, which comprises approximately 1.4M dialogues produced using OpenAI’s Turbo APIs. From UltraChat200k, We randomly sample $50$k prompts and use the base model to generate the synthetic responses. We subsequently follow the optimization method described in Section 4.1 for further training. In multiple iterations, we leverage the synthetic data from the most recent iteration and add to the newly generated synthetic data, therefore resulting in a synthetic dataset size of $50$k at iteration 00 and $100$k at iteration $1$, $2$ and $3$. At each iteration, we train our model for $2$ epochs.
@@
Evaluation. We employed the widely used Huggingface Open LLM Leaderboard (Beeching et al., 2023) as our evaluation benchmark, using the same Language Model Evaluation Harness library (Gao et al., 2023b). This leaderboard encompasses 6 different datasets, each focusing on a a specific capability of LLMs. Collectively, these datasets provide a thorough assessment framework, evaluating LLMs on commonsense reasoning (Arc (Clark et al., 2018), HellaSwag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2021)), multi-task language understanding (MMLU(Hendrycks et al., 2020)), human falsehood mimic (TruthfulQA (Lin et al., 2021)) and math problem solving (GSM8k (Cobbe et al., 2021)). We leave further implementation details to Appendix B with detailed evaluation setting adopted by both the leaderboard and our experiments.
6.2 SPIN Effectively Improves Benchmark Performance
In Figure 2, we demonstrate the effectiveness of SPIN using HuggingFace Open LLM Leaderboard as the evaluation. In Figure 3, we compare the performance of our fine-tuned model by SPIN after iterations 0 to 3 with the base model zephyr-7b-sft-full on each task included in the leaderboard. Detailed performances are presented in Table 4 in Appendix B. We can observe that SPIN exhibits remarkable effectiveness in improving the model’s performance by further leveraging the SFT dataset, on which the base model has already been fully fine-tuned. At iteration 00, where model responses are generated from zephyr-7b-sft-full, we observe an overall improvement of $2.66\%$ on the average score. The improvement is particularly significant on the TruthfulQA and GSM8k benchmarks, with improvement exceeding $5\%$ and $10\%$ respectively. At iteration $1$, we employ the LLM model from iteration 00 to generate new responses for SPIN, adhering to the procedure outlined in Algorithm 1. This iteration yields further enhancements of $1.32\%$ on average, and especially significant on the Arc Challenge and TruthfulQA benchmarks. Subsequent iterations continue this trend of incremental improvement across various tasks. Meanwhile, the improvement at iteration $t+1$ is naturally smaller than that at iteration $t$. As the iterative training progresses, the degree of improvement gradually approaches zero, suggesting that the model has reached a limiting point in the last iteration.
Comparison with DPO. zephyr-7b-beta is a model derived from zephyr-7b-sft-full, trained with DPO on approximately $62$k preference data. This data, the UltraFeedback Binarized dataset (Cui et al., 2023) [^2] https://huggingface.co/datasets/HuggingFaceH4/ultrafeedback_binarized, comprises both chosen and rejected completions evaluated by GPT-4. We note that, DPO requires either human input or advanced language model feedback to determine the preference, making data generation a rather expensive procedure. In contrast, our SPIN only requires the initial model itself. Moreover, unlike DPO which requires new data source, our method exclusively leverages the existing SFT dataset. In Figure 3, we show the performance comparison of SPIN at iterations 0 and 1 (employing $50$k SFT data) with DPO training, from the same SFT checkpoint. We can observe that, while DPO leverages more data from new sources, SPIN based on the existing SFT data can already achieve comparable average performance to DPO training at iteration 0. From iteration 1, SPIN even surpasses the performance of DPO on the leaderboard benchmark.
(a) Arc Challenge.
(b) TruthfulQA.
(c) Average.
6.3 Ablation Studies
In this subsection, we examine the effect of synthetic dataset size and training epochs within an iteration. Our analysis demonstrates the effectiveness of the synthetic data used by SPIN compared to the SFT data, as well as the necessity of iterative training in SPIN. In Appendix B, we present assessment of SPIN on additional benchmark tasks.
Training Size. We investigate the effect of varying training data size on the performance of SPIN. In Figure 5, we demonstrate the effect of training size for SPIN during iteration 00 and additionally compare with SFT with the full original dataset. Specifically, for the SFT baseline, we fully fine-tune Mistral-7B on Ultrachat200k for three epochs and report first epoch performance as the starting point (with x-axis 0) in the figure for SFT. For SPIN, we report the zephyr-7b-sft-full checkpoint as the starting point, which has also been fine-tuned on Ultrachat200k for one epoch. We select the training size of SPIN at iteration 0 to be 14k, 26k, and 50k and generate the data accordingly, ensuring that the larger dataset encompasses the smaller dataset. The performance of SPIN was then evaluated after 1 epoch of self-play fine-tuning for each training size. We can observe that, while SPIN results in notable improvement with increasing training sizes, SFT on further epochs 2 and 3 fails to yield more than $1\%$ improvement. Additional results are deferred to Appendix B.
Iterative Training v.s. Training for More Epochs. We further study the training within iteration 00 and compare with the performance achieved in iteration $1$, particularly contrasting the test performance obtained from extended training duration with that from next iteration. Figure 4 depicts the performance trajectory of the model trained using SPIN over multiple epochs at iteration 0. It is evident that the most substantial improvement occurs during the first two epochs, followed by only modest gains in subsequent epochs. Notably, SPIN exhibits robustness and stability; extending the training duration does not diminish performance but rather maintains a rather consistent level. Nevertheless, the observation suggests an inherent limitation to the performance achievable within a single iteration, thereby underscoring the necessity for iterative training. As shown by the test performance achieved at iteration 1 in the figures, extending the training in iteration 0 fails to reach the performance comparable to iteration 1.
7 Conclusion and Discussion
Limitation and Future Work. Our theoretical results demonstrate that the optimization process of SPIN converges if and only if the LLM’s distribution aligns with $p_{\mathrm{data}}$. Therefore, our study focuses on a fixed target data distribution generated by humans, which inherently imposes a ceiling on the performance of fine-tuned LLM. Exploring the dynamically changing target data distribution is an important direction to overcome this limitation and elevate the LLM’s performance beyond this ceiling or even to a super-human level. Moreover, considering the resource demands of synthetic data generation, another promising avenue for further exploration is to reduce the volume of required synthetic data.
Acknowledgement
We thank the anonymous reviewers and area chair for their helpful comments. ZC, YD, HY, KJ, and QG are supported in part by the National Science Foundation CAREER Award 1906169, IIS-2008981, CHE-2247426 and the Sloan Research Fellowship. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing any funding agencies.
Impact Statement
This paper presents work whose goal is to advance the field of Large Language Models. We believe that our work contribute meaningfully to the field, specifically on leveraging synthetic data to enhance LLM without the requirement for human preference annotations. The synthetic data generated by SPIN may be used to further augment the training of various language models. Moreover, SPIN demonstrated a substantial improvement in LLMs’ capabilities, opening new avenues for their application in various downstream tasks. This advancement underscores the transformative potential of LLM fine-tuning in both technological and societal contexts.
Appendix A Further Related Work
Curriculum Learning. In deep learning, it has been observed that training models using data samples arranged in a strategically meaningful order can lead to improved performance compared to training on randomly shuffled data. This approach is commonly known as curriculum learning (Bengio et al., 2009; Soviany et al., 2022). Initial studies in curriculum learning introduced efficient algorithms that adhere to an ‘easy-to-hard’ progression (Spitkovsky et al., 2009; Kumar et al., 2010; Lee & Grauman, 2011; Zhang et al., 2015). In the field of Natural Language Processing (NLP), criteria such as sentence length and term frequency are commonly utilized (Cirik et al., 2016; Zhang et al., 2018; Liu et al., 2018). More recent developments include the application of curriculum learning algorithms in multi-modal learning (Liu et al., 2021; Wu et al., 2022). Our work shares a similar idea to curriculum learning, wherein the training data evolves iteratively—beginning with responses that are easy to distinguish from human-annotated data and gradually progressing to more challenging instances.
Generative Adversarial Networks. Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) represent a distinct class of generative models, characterized by their unique adversarial process. To enhance training stability and data quality, Mao et al. (2017) introduced the Least Squares GAN, employing a least squares loss function for the discriminator. A significant advancement in GANs involves the use of Integral Probability Metrics (IPM) (Müller, 1997), particularly highlighted in the development of Wasserstein GAN by Arjovsky et al. (2017). This model employs IPM in its loss design, enhancing training stability. Since then, IPMs have become crucial in GAN design (Mroueh & Sercu, 2017; Gulrajani et al., 2017), particularly in constraining the discriminator to a specific function class, thereby preventing it from overpowering the generator. Furthermore, Jolicoeur-Martineau (2018) generalized IPM-based GANs by introducing relativistic discriminator and proposed Relativistic GAN. It is worth noting that the objective function defined in our (4.2) is similar to Relativistic GAN (Jolicoeur-Martineau, 2018) and reduces to an IPM framework in Wasserstein GAN (Arjovsky et al., 2017) with a linear loss. However, our approach differs in both the choice of the function class and the training procedure. Inspired by GAN, Cheng et al. (2023) proposed an adversarial learning framework named Adversarial Preference Optimization (APO) that trains the LLM and a reward model in an adversarial game. Similarly related to the spirit of our method, Generative Adversarial Imitation Learning (GAIL) (Ho & Ermon, 2016) was proposed to train separate discriminator and policy networks in each iteration. In contrast to the above methods, SPIN relies on self-play where both the main player and the opponent player are the same LLM from two consecutive iterations.
Alignment with AI Feedback. The objective of alignment is to fine-tune LLMs to align with human intentions. In addition to using human demonstrations, AI feedback is emerging as a crucial component in the alignment process. Constitutional AI (Bai et al., 2022b) leveraged AI feedback to align language models through a combination of both supervised learning and reinforcement learning (RL) phases. In the RL phase, (Bai et al., 2022b) applied Reinforcement Learning from AI Feedback (RLAIF), training a reward model using AI-generated preferences, followed by RL using the reward. Lee et al. (2023) demonstrated that AI feedback can achieve comparable or superior performance to human feedback in RL fine-tuning. They also demonstrated that RLAIF can improve upon an SFT policy even when the LLM labeler has the same size as the policy. Saunders et al. (2022) studied the scaling properties of self-critique and introduced a framework for comparing the critique ability to generation and discrimination ability. Self-critique models employ the LLM itself to generate natural language critiques through behavioral cloning, assisting human evaluators.
Appendix B Experiments
B.1 Hyperparameters and Implementation Details
@@
[…]
B.2 Training Overhead
The cost overhead introduced by SPIN is mostly the generation of synthetic data from the LLM that we train. The cost of the fine-tuning process remains computationally equal to that of SFT and DPO. We report both the generation and training time for SPIN in Table 2 . Results were obtained using a machine with 8xA100 (80G) GPUs. For per 64 examples, the generation time and training time are 6.69s and 10s respectively.
Table 2: Generation and Training Times for Different Iterations
Iteration Iter 0 Iter 1 Iter 2 Iter 3
Process Generation Training Generation Training Generation Training Generation Training
Time 1.45h 4.32h 1.45h 8.64h 1.45h 8.64h 1.45h 8.64h
It is evident that the generation time is dominated by the training time at each iteration. The estimated time in Table 2 is based on the fact that we generate 50k examples per iteration. Please note that the doubled training time from iter 1 to iter 3 is attributed to the utilization of a double-sized dataset (the combination of 50k synthetic data from the previous iteration and 50k synthetic data in the current iteration), as discussed in our Section 6.1.
B.3 Additional Experiment Result for SPIN+DPO
SPIN requires only the SFT data to improve over the traditional SFT stage and can sit between SFT and RL finetuning. Suppose additional preference data is provided, we can use the additional data to further improve the performance of the model after SPIN using RL fine-tuning.
Starting at SPIN iteration 3, we further train the model with DPO for two epochs on the $62$k preference data from the UltraFeedback Binarized dataset (Cui et al., 2023), which consists of both chosen and rejected responses evaluated by GPT-4. Detailed performances are presented in Table 3.
Table 3: Performance of SPIN + DPO based on zephyr-7b-sft-full across HuggingFace Open LLM Leaderboard datasets, compared with all baselines. We also denote the average improvement over last iteration in the Average column.
Model Arc TruthfulQA Winogrande GSM8k HellaSwag MMLU Average
zephyr-7b-dpo-full 63.65 55.19 72.61 33.43 84.44 58.52 61.31
zephyr-7b-sft-full 60.41 43.73 74.19 26.76 82.85 60.92 58.14
SPIN iteration 00 63.40 49.18 72.69 35.10 84.38 60.03 $60.80_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}(+2.66)}}$
SPIN iteration $1$ 65.19 55.17 72.30 35.78 84.96 59.34 $62.12_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}(+1.32)}}$
SPIN iteration $2$ 65.96 54.91 73.56 38.06 85.41 59.93 $62.97_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}(+0.85)}}$
SPIN iteration $3$ 65.87 54.90 73.72 38.97 85.54 59.99 $63.16_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}(+0.19)}}$
SPIN iteration $3$ + DPO 66.47 60.07 78.06 37.98 86.17 59.68 $64.05_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}(+0.89)}}$
We can observe that the checkpoint trained by SPIN can be further improved using DPO, yielding an enhancement of $0.89\%$ on average. Notably, the improvement is particularly significant on the TruthfulQA benchmark with around $5\%$.
B.4 Further Experiment Results
In Table 4, we show the detailed performance of SPIN at different iterations on each of the task in Open LLM Leaderboard. In Table 5, we also show the performance of SFT from zephyr-7b-sft-full on Ultrachat200k for one epoch. While self-play fine-tuning with synthetic data from zephyr-7b-sft-full effectively improves its performance, simply fine-tuning it again on the SFT data leads to degraded performance, as similarly observed in Figure 5.
@@ B.4 Further Experiment Results
SPIN iteration 00 63.40 49.18 72.69 35.10 84.38 60.03 $60.80_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}(+2.66)}}$
SPIN iteration $1$ 65.19 55.17 72.30 35.78 84.96 59.34 $62.12_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}(+1.32)}}$
SPIN iteration $2$ 65.96 54.91 73.56 38.06 85.41 59.93 $62.97_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}(+0.85)}}$
SPIN iteration $3$ 65.87 54.90 73.72 38.97 85.54 59.99 $63.16_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}(+0.19)}}$
@@
Further Investigation on More Tasks. Here, we further investigate the performance of SPIN on a broader variety of tasks, including MT-Bench (Zheng et al., 2023), Big-Bench (bench authors, 2023) and OpenBookQA (Mihaylov et al., 2018) in addition to the Open LLM Leaderboard tasks. Specifically, we use the following tasks from Big-Bench-Hard for a more comprehensive evaluation, including Causal Judgment (causal reasoning), Sports Understanding (commonsense reasoning) and Formal Fallacies (logical reasoning). In Table 6, we show the resulting scores of SPIN on MT-Bench as well as those tasks from Big-Bench. In Figure 6, we detail the model performances on MT-Bench with regard to different types of questions. We can see a notably robust improvement in the performance of SPIN on various tasks besides the HuggingFace Benchmark, without major degradation. Notably, on MT-Bench, the model fine-tuned by SPIN has surpassed the performance of vicuna-13b-v1.5 (Chiang et al., 2023) with a score of $6.57$.
@@
SPIN iteration 00 $6.46_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}(+0.52)}}$ 57.75 51.6 95.2 46.8
SPIN iteration $1$ $6.65_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}(+0.19)}}$ 58.82 51.2 95.2 47.2
SPIN iteration $2$ $6.78_{{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}(+0.13)}}$ 59.36 51.2 94.4 47.6
@@
B.5 Generation Examples
In Tables 7 and 8, we further provide the generation examples of our fine-tuned model by SPIN from different iterations. We can observe an improvement in response quality as compared to the generation of the SFT checkpoint. Meanwhile, the model generations at higher iterations typically becomes more concise than iteration 00 and resemble the ground truth completion better.
Prompt How does the economic growth of Horsham compare to nearby towns?
Ground truth I do not have access to the latest economic data on horsham and nearby towns. However, you can check the latest economic statistics and reports provided by the local government or the official websites of horsham and nearby towns for a detailed comparison of their economic growth.